a boy is running along circular track of 15 m diameter. if in 20 seconds he covers half the track, then find his angular speed and linear speed?
1
Expert's answer
2019-09-18T09:47:00-0400

The period of a boy motion along circular track "T=2t=40\:\rm s."

The angular speed


"\omega=\frac{2\pi}{T}=\frac{2\pi}{40\:\rm s}=\frac{\pi}{20}\:\rm s^{-1}=0.079\:s^{-1}."

The linear speed


"v=\frac{2\pi R}{T}=\frac{2\pi\times 7.5\:\rm m}{40\:\rm s}=\frac{3\pi}{8}\:\rm m/s=1.178\: m/s."